OpenStudy (zeesbrat3):

A cylinder labeled X has a height of 10 inches and a radius of 15 inches. Another cylinder labeled Y has a height of 40 inches and a radius of 60 inches. The volume of cylinder Y is _______ times the volume of cylinder X. I think it is 64 times bigger, but Im not sure.

ganeshie8 (ganeshie8):

thats correct !

ganeshie8 (ganeshie8):

you just compute the volumes for both. and take ratio

OpenStudy (zeesbrat3):

Thanks. I just wasnt sure if you had to square root it or not

ganeshie8 (ganeshie8):

hmm what ?

OpenStudy (zeesbrat3):

Well 64 can be square rooted, and I know that with certain problems, you have to. Unless Im probably thinking to hard on this problem and I got myself all confused, which is likely.

ganeshie8 (ganeshie8):

you thinking too hard. and you shouldnt square root 64, becos :

ganeshie8 (ganeshie8):

question is asking how many times the volume of Y is bigger than volume of X

ganeshie8 (ganeshie8):

Volume of Y = \(\pi. r^2. h = \pi . 60^2 . 40\) Volume of Y = \(\pi. r^2. h = \pi . 15^2 . 10\) you take Y/X it gives how many times Y is bigger than X

OpenStudy (zeesbrat3):

Okay, I guess I got used to harder problems. Thanks. :)

ganeshie8 (ganeshie8):

sorry second line is "Volume of X"

ganeshie8 (ganeshie8):

haha happens with me...

OpenStudy (zeesbrat3):

but your too smart to make simple mistakes!!

ganeshie8 (ganeshie8):

oh.. ty :)